Washington Air Company is considering the purchase of a helicopter for connecting services between the​ company's base airport and the new intercounty airport being built about 30 miles away. It is believed that the chopper will be needed only for six years until the Rapid Transit Connection is phased in. The estimates on two types of helicopters under​ consideration, the Whirl 2B and the ROT​ 8, are given in the table below. Assuming that the Whirl 2B will be available in the future with identical​ costs, what is the annual cost advantage of selecting the ROT​ 8? (Use an interest rate of 14​%.)

First cost Annual maintenance Salvage value Useful life in years The Whirl 2B $92,000 S3,500 $15,000 The ROT 8 $109,000 $9,500 $20,000

The annual cost advantage of selecting the ROT 8 is ​$_________.

​(Round to the nearest​ dollar.)
Answer #1

Annual worth (AW) of costs is computed as follows.

AW, Whirl 2B ($) = 92,000 x A/P(14%, 3) + 3,500 - 15,000 x P/F(14%, 3) x A/P(14%, 3)

= 92,000 x 0.4307** + 3,500 - 15,000 x 0.6750** x 0.4307**

= 39,624.4 + 3,500 - 4,360.84

= 38,763.56

AW, ROT 8 ($) = 109,000 x A/P(14%, 6) + 9,500 - 20,000 x P/F(14%, 6) x A/P(14%, 6)

= 109,000 x 0.2572** + 9,500 - 20,000 x 0.4556** x 0.2572**

= 28,034.8 + 9,500 - 2,343.61

= 35,191.19

Annual cost advantage of selecting ROT 8 ($) = 38,763.56 - 35,191.19 = 3,572.37 ~ 3,572

**From A/P and P/F Factor tables